OpenStudy (conqueror):

Lines A and B form a linear system. Line A has slope of 3 and an x-intercept of (9, 0). Line B has a slope of 3 and has an x-intercept of (9, 0). How are these lines characterized? parallel intersecting coincidental not enough information

OpenStudy (texaschic101):

I started working this out by using the point slope formula y - y1 = m(x - x1) slope(m) = 3 (9,0)..x1 = 9 and y1 = 0 now we sub y - 0 = 3(x - 9) y = 3x - 18 -3x + y = -18 3x - y = 18 both of your lines will be the same, therefore, your lines are coincidental.